# Palm distributions of superposed point processes for statistical inference

### v1

@@ Acknowledgments
$j_{k}(x_{1},\ldots,x_{k})=k!\mathsf{P}(N=k)$
$\quad\times\sum_{\bm{t}\in\mathcal{T}_{k}}\left[\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}}\prod_{j:t_{j}=h}\kappa(x_{i};\theta)G_{0}(\mathrm{d}\theta)\right]\frac{\exp\left\{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)\right\}\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma)}{k!\mathsf{P}(N=k)},$
where the fraction term corresponds to $\mathsf{P}(\bm{T}=\bm{t})$, where the law of $\bm{T}$ is defined in the statement. In conclusion, the thesis follows by expressing the last equation as an expected value with respect to $\bm{T}$. ∎
S4.4 Expectation-maximization algorithm for maximum likelihood estimation
Let $\Phi\sim\textsc{sncp}(\kappa,\nu)$ and denote by $\psi$ the set of parameters of $\Phi$, which enters the definition of $\kappa$ and $\nu$. Define the likelihood $L(\psi;\bm{x})=j_{k}(x_{1},\ldots,x_{k})$ as in Theorem 3, for $\bm{x}=(x_{1},\ldots,x_{k})$. The peculiar structure of the likelihood naturally motivates the use of an em algorithm for its maximization, which we develop in the following. Let us introduce the following shorthand notation
$\begin{split}p(\bm{x}\,|\,\bm{t},\psi)=\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)G_{0}(\mathrm{d}\theta),&\quad p(\bm{t}\,|\,\psi)\propto\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma),\\ c(\psi)&=k!\,\mathsf{P}(\Phi(\mathbb{X})=k),\end{split}$ (S12)
where $\bm{t}=(t_{1},\ldots,t_{k})\in\mathcal{T}_{k}$. Then
$L(\psi;\bm{x})=c(\psi)p(\bm{x}\,|\,\psi)=c(\psi)\sum_{\bm{t}\in\mathcal{T}_{k}}p(\bm{x},\bm{t}\,|\,\psi)=c(\psi)\sum_{\bm{t}\in\mathcal{T}_{k}}p(\bm{x}\,|\,\bm{t},\psi)p(\bm{t}\,|\,\psi).$ (S13)
Following em_notes, let
$Q(\psi;\psi_{old}):=\mathsf{E}\left[\ln\left\{c(\psi)p(\bm{x},\bm{T}\,|\,\psi)\right\}\,|\,\bm{x},\psi_{old}\right].$ (S14)
The em algorithm alternates computing $Q(\psi;\psi_{old})$ for the current values of parameters $\psi_{old}$ (E-step) and finding a new value $\psi_{new}$ maximizing $Q(\psi;\psi_{old})$ (M-step). See em_notes for further details and justifications. Since the expecation in (S14) is not available analytically, we employ a stochastic approximation of the E-step as in wei1990monte and approximate
$Q(\psi;\psi_{old})\approx\frac{1}{U}\sum_{u=1}^{U}\ln\left\{c(\psi)p(\bm{x},\bm{T}^{(u)}\,|\,\psi)\right\},\qquad\bm{T}^{(1)},\ldots,\bm{T}^{(U)}\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}p(\cdot\,|\,\bm{x},\psi_{old}),$
In our examples, we find that $U=1$ gives satisfactory results. To sample $\bm{T}$ from $p(\cdot\,|\,\bm{x},\psi_{old})$, we employ a standard Gibbs sampler developed for Bayesian nonparametric mixture models, see Algorithm 3 in neal2000markov. The Gibbs sampler starts from an initial configuration for $\bm{T}$ and updates it through a number of iterations; at each iteration, $\bm{T}$ is updated by sequentially updating each of its elements $T_{j}$, as $j=1,\ldots,k$. In particular, for $\ell=1,\ldots,k$, the update of $T_{\ell}$ given the current values of the remaining latent indicators $\bm{T}_{-\ell}=(T_{j}:j\neq\ell)$ follows from its full-conditional distribution. Let $C$ be the number of clusters in the partition induced by $\bm{T}_{-\ell}$, with cluster cardinalities $n_{h}$ and associated distinct label $t^{*}_{h}$, as $h=1,\ldots,C$. The update for $T_{\ell}$, denoted with $T_{\ell}^{new}$, follows
$p(T_{\ell}^{new}\,|\,\bm{T}_{-\ell},\bm{x})\propto\sum_{h=1}^{C}p(\bm{x}\,|\,\bm{t}^{\prime}_{h})p(\bm{t}^{\prime}_{h})\delta_{t^{*}_{h}}+p(\bm{x}\,|\,\bm{t}^{\prime}_{new})p(\bm{t}^{\prime}_{new})\delta_{\min\{i:i\neq t^{*}_{h},h=1,\ldots,C\}},$
where $\bm{t}^{\prime}_{h}$ is obtained by inserting the value $t^{*}_{h}$ in the $\ell$-th position of the vector $\bm{T}_{-\ell}$, shifting subsequent elements one place to the right. Similarly, $\bm{t}^{\prime}_{new}$ is obtained by inserting the value $\min\{i:i\neq t^{*}_{h},h=1,\ldots,C\}$ in the $\ell$-th position of the vector $\bm{T}_{-\ell}$. Note that the number of clusters in the partition induced by $\bm{t}^{\prime}_{h}$ is $C$, for any $h$, while the number of clusters in the partition induced by $\bm{t}^{\prime}_{new}$ is $C+1$. The probability density functions $p(\bm{x}\,|\,\bm{t})$ and $p(\bm{t})$ are described in (S12). In our experiments, we run the Gibbs sampler for 20 iterations and keep the last partition visited as a draw from $p(\cdot\,|\,\bm{x},\psi_{old})$. The M-step is carried out by a simple gradient descent method, obtaining the gradients of $Q$ via automatic differentiation.
S4.5 Proofs of the auxiliary results in Appendix S4.1
Proof of Lemma S1
@@
$\mathcal{L}_{\Phi}(f)$ $=\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi(\mathrm{d}x)\right\}\right]=\mathsf{E}\left[\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi(\mathrm{d}x)\right\}\,|\,\Lambda\right]\right]$
$=\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}(1-e^{-f(x)})\sum_{j\geq 1}\gamma_{j}\kappa(x;\theta_{j})\mathrm{d}x\right\}\right]$
$=\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\int_{\mathbb{X}}(1-e^{-f(x)})\kappa(x;\theta)\mathrm{d}x\,\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\right]$
@@
Proof of Lemma S2
Point (i). For the Laplace functional of $\Phi_{\zeta_{\bm{x}}}$, with $\bm{x}=(x_{1},\ldots,x_{k})$, the proof follows from direct computation. Let $f:\mathbb{X}\to\mathbb{R}_{+}$ be any measurable function, then
$\mathcal{L}_{\Phi_{\zeta_{\bm{x}}}}(f)$ $=\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi_{\zeta_{\bm{x}}}(\mathrm{d}x)\right\}\right]=\mathsf{E}\left[\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi_{\zeta_{\bm{x}}}(\mathrm{d}x)\right\}\,|\,\zeta_{\bm{x}}\right]\right]$
$=\mathsf{E}\left[\exp\left\{-\int_{\mathbb{X}}(1-e^{-f(x)})\gamma_{\bm{x}}\kappa(x;\theta_{\bm{x}})\mathrm{d}x\right\}\right]$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\gamma\int_{\mathbb{X}}(1-\exp\{-f(x)\})\kappa(x;\theta)\mathrm{d}x\right\}\gamma^{k}\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)/\eta(\bm{x}).$
The third equality follows from the fact that $\Phi_{\zeta_{\bm{x}}}\,|\,\zeta_{\bm{x}}$ is distributed as Poisson process.
Point (ii). Still by direct computation, the mean measure of $\Phi_{\zeta_{\bm{x}}}$ writes as
$M_{\Phi_{\zeta_{\bm{x}}}}(\mathrm{d}y)$ $=\mathsf{E}\left[\mathsf{E}\left[\Phi_{\zeta_{\bm{x}}}(\mathrm{d}y)\,|\,\zeta_{\bm{x}}\right]\right]=\mathsf{E}\left[\gamma_{\bm{x}}\kappa(y;\theta_{\bm{x}})\mathrm{d}y\right]=$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(y;\theta)\mathrm{d}y\;\gamma^{k}\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)/\eta(\bm{x})$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{k+1}\kappa(y;\theta)\;\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}y/\eta(\bm{x})$
$=\eta(\bm{x},y)\mathrm{d}y/\eta(\bm{x})$
@@
Point (iii). By (BaBlaKa, Proposition 3.2.5), if $\Psi$ is a Cox process directed by the random measure $\mu$, then the reduced Palm version $\Psi^{!}_{y}$ is a Cox process directed by $\mu_{y}$ (the Palm version of $\mu$ at $y$). Now, observe that $\Phi_{\zeta_{\bm{x}}}$ is a Cox process directed by the random measure $\mu(\mathrm{d}x)=\gamma_{{\bm{x}}}\kappa(x;\theta_{{\bm{x}}})\,\mathrm{d}x$, with $(\theta_{{\bm{x}}},\gamma_{{\bm{x}}})\sim f_{{\bm{x}}}(\mathrm{d}\theta\,\mathrm{d}\gamma)$ and $f_{{\bm{x}}}$ is defined in the statement of the present lemma. Consequently, the reduced Palm version $\left(\Phi_{\zeta_{{\bm{x}}}}\right)^{!}_{y}$ is a Cox process directed by $\mu_{y}$. We are then left with determining the distribution of $\mu_{y}$, which we derive by computing its Laplace functional $\mathcal{L}_{\mu_{y}}(f)$, for any measurable function $f:\mathbb{X}\to\mathbb{R}_{+}$. To this end, rely on (BaBlaKa, Proposition 3.2.1), which states that, for any measurable functions $f,g:\mathbb{X}\to\mathbb{R}_{+}$, the Palm distribution of a random measure $\mu$ satisfies
$evalat{\frac{\partial}{\partial t}\mathcal{L}_{\mu}(f+tg)}{t=0}=-\mathsf{E}\left[\mu(g)e^{-\mu(f)}\right]=-\int_{\mathbb{X}}g(y)\mathcal{L}_{\mu_{y}}(f)M_{\mu}(\mathrm{d}y),$ (S15)
where $\mu(f):=\int_{\mathbb{X}}f(x)\mu(\mathrm{d}x)$. Therefore, since $\mu(\mathrm{d}x)=\gamma_{{\bm{x}}}\kappa(x;\theta_{{\bm{x}}})\,\mathrm{d}x$, the term $\mathsf{E}\left[\mu(g)e^{-\mu(f)}\right]$ writes as
$\mathsf{E}\left[\mu(g)e^{-\mu(f)}\right]$ $=\mathsf{E}\left[\int_{\mathbb{X}}g(y)\gamma_{\bm{x}}\kappa(y;\theta_{\bm{x}})\mathrm{d}y\cdot\exp\left\{-\int_{\mathbb{X}}f(z)\gamma_{\bm{x}}\kappa(z;\theta_{\bm{x}})\mathrm{d}z\right\}\right]$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\int_{\mathbb{X}}g(y)\gamma\kappa(y;\theta)\mathrm{d}y\cdot\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\cdot f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)$
$=\int_{\mathbb{X}}g(y)\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\gamma\kappa(y;\theta)f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\cdot\mathrm{d}y.$ (S16)
Observing that $M_{\mu}(\mathrm{d}y)=\mathsf{E}[\gamma_{\bm{x}}\kappa(y;\theta_{\bm{x}})\mathrm{d}y]=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(y;\theta)f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\,\mathrm{d}y$, (S16) can be expressed as
$\mathsf{E}\left[\mu(g)e^{-\mu(f)}\right]=\int_{\mathbb{X}}g(y)\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\frac{\gamma\kappa(y;\theta)f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)}{\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{\prime}\kappa(y;\theta^{\prime})f_{\bm{x}}(\mathrm{d}\theta^{\prime}\,\mathrm{d}\gamma^{\prime})}\cdot M_{\mu}(\mathrm{d}y)$
@@
$\mathcal{L}_{\mu_{y}}(f)=\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\frac{\gamma\kappa(y;\theta)f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)}{\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{\prime}\kappa(y;\theta^{\prime})f_{\bm{x}}(\mathrm{d}\theta^{\prime}\,\mathrm{d}\gamma^{\prime})},$
@@
$(\theta^{*},\gamma^{*})\sim f_{*}(\mathrm{d}\theta\,\mathrm{d}\gamma)\propto\gamma\kappa(y;\theta)f_{\bm{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\propto f_{(\bm{x},y)}(\mathrm{d}\theta\,\mathrm{d}\gamma).$
@@
$\left(\Phi_{\zeta_{{\bm{x}}}}\right)^{!}_{y}\,|\,(\theta^{*},\gamma^{*})\sim\mathrm{PP}(\gamma^{*}\kappa(x;\theta^{*})\,\mathrm{d}x),\qquad(\theta^{*},\gamma^{*})\sim f_{(\bm{x},y)}(\mathrm{d}\theta\,\mathrm{d}\gamma),$
which is equivalent to say $\left(\Phi_{\zeta_{{\bm{x}}}}\right)^{!}_{y}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\Phi_{\zeta_{({\bm{x}},y)}}$, and the thesis is proved.
Proof of Lemma S3
@@
$M^{(k)}_{\Phi}(B_{1},\ldots,B_{k})$ $=\mathsf{E}\left[\sum_{X_{1},\ldots,X_{k}\in\Phi}^{\neq}\prod_{j=1}^{k}\mathbbm{1}_{B_{j}}(X_{j})\right]=\mathsf{E}\left[\mathsf{E}\left[\sum_{X_{1},\ldots,X_{k}\in\Phi}^{\neq}\prod_{j=1}^{k}\mathbbm{1}_{B_{j}}(X_{j})\,|\,\Lambda\right]\right]$
$=\mathsf{E}\left[M^{(k)}_{\Phi\,|\,\Lambda}(B_{1},\ldots,B_{k})\right]=\mathsf{E}\left[\prod_{j=1}^{k}M_{\Phi\,|\,\Lambda}(B_{j})\right]$
$=\mathsf{E}\left[\prod_{j=1}^{k}\int_{B_{j}}\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(x_{j};\theta)\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}x_{j}\right]$
$=\mathsf{E}\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{k}}\prod_{j=1}^{k}\gamma_{j}\int_{B_{j}}\kappa(x_{j};\theta_{j})\mathrm{d}x_{j}\Lambda^{k}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma})\right]$
$=\int_{(\mathbb{X}\times\mathbb{R}_{+})^{k}}\prod_{j=1}^{k}\gamma_{j}\int_{B_{j}}\kappa(x_{j};\theta_{j})\mathrm{d}x_{j}M_{\Lambda}^{k}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma}),$ (S17)
@@
Moreover, from (BaBlaKa, Lemma 14.E.4), the following holds true for any point process $\xi$ on a Polish space $\mathbb{Y}$ (embedded with the Borel $\sigma$-algebra $\mathscr{Y}$), for any $A_{1},\ldots,A_{k}\in\mathscr{Y}$,
@@
where $k_{h}$ is the cardinality of set $J_{h}$. For $\bm{x}_{J_{h}}=(x_{m}:m\in J_{h})$, define
$\eta(\bm{x}_{J_{h}})=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{k_{h}}\prod_{m\in J_{h}}\kappa(x_{m};\theta)M_{\Lambda}(\mathrm{d}\theta\,\mathrm{d}\gamma).$

### v2

@@ Acknowledgments
$j_{k}(x_{1},\ldots,x_{k})=k!{\rm pr}(N=k)$
$\quad\times\sum_{\underline{t}\in\mathcal{T}_{k}}\left[\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{X}}\prod_{j:t_{j}=h}\kappa(x_{i};\theta)G_{0}(\mathrm{d}\theta)\right]\frac{\exp\left\{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)\right\}\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma)}{k!{\rm pr}(N=k)},$
where the fraction term corresponds to ${\rm pr}(\underline{T}=\underline{t})$, where the law of $\underline{T}$ is defined in the statement. In conclusion, the thesis follows by expressing the last equation as an expected value with respect to $\underline{T}$. ∎
S6.4 Proofs of the auxiliary results in Appendix S6.1
Proof of Lemma˜S1
@@
$\mathcal{L}_{\Phi}(f)$ $=E\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi(\mathrm{d}x)\right\}\right]=E\left[E\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi(\mathrm{d}x)\right\}\,|\,\Lambda\right]\right]$
$=E\left[\exp\left\{-\int_{\mathbb{X}}(1-e^{-f(x)})\sum_{j\geq 1}\gamma_{j}\kappa(x;\theta_{j})\mathrm{d}x\right\}\right]$
$=E\left[\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\int_{\mathbb{X}}(1-e^{-f(x)})\kappa(x;\theta)\mathrm{d}x\,\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\right]$
@@
Proof of Lemma˜S2
Point (i). For the Laplace functional of $\Phi_{\zeta_{\underline{x}}}$, with $\underline{x}=(x_{1},\ldots,x_{k})$, the proof follows from direct computation. Let $f:\mathbb{X}\to\mathbb{R}_{+}$ be any measurable function, then
$\mathcal{L}_{\Phi_{\zeta_{\underline{x}}}}(f)$ $=E\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi_{\zeta_{\underline{x}}}(\mathrm{d}x)\right\}\right]=E\left[E\left[\exp\left\{-\int_{\mathbb{X}}f(x)\Phi_{\zeta_{\underline{x}}}(\mathrm{d}x)\right\}\,|\,\zeta_{\underline{x}}\right]\right]$
$=E\left[\exp\left\{-\int_{\mathbb{X}}(1-e^{-f(x)})\gamma_{\underline{x}}\kappa(x;\theta_{\underline{x}})\mathrm{d}x\right\}\right]$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\gamma\int_{\mathbb{X}}(1-\exp\{-f(x)\})\kappa(x;\theta)\mathrm{d}x\right\}\gamma^{k}\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)/\eta(\underline{x}).$
The third equality follows from the fact that $\Phi_{\zeta_{\underline{x}}}\,|\,\zeta_{\underline{x}}$ is distributed as Poisson process.
Point (ii). Still by direct computation, the mean measure of $\Phi_{\zeta_{\underline{x}}}$ writes as
$M_{\Phi_{\zeta_{\underline{x}}}}(\mathrm{d}y)$ $=E\left[E\left[\Phi_{\zeta_{\underline{x}}}(\mathrm{d}y)\,|\,\zeta_{\underline{x}}\right]\right]=E\left[\gamma_{\underline{x}}\kappa(y;\theta_{\underline{x}})\mathrm{d}y\right]=$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(y;\theta)\mathrm{d}y\;\gamma^{k}\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)/\eta(\underline{x})$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{k+1}\kappa(y;\theta)\;\prod_{j=1}^{k}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}y/\eta(\underline{x})$
$=\eta(\underline{x},y)\mathrm{d}y/\eta(\underline{x})$
@@
Point (iii). By (Baccelli et al., 2020, Proposition 3.2.5), if $\Psi$ is a Cox process directed by the random measure $\mu$, then the reduced Palm version $\Psi^{!}_{y}$ is a Cox process directed by $\mu_{y}$ (the Palm version of $\mu$ at $y$). Now, observe that $\Phi_{\zeta_{\underline{x}}}$ is a Cox process directed by the random measure $\mu(\mathrm{d}x)=\gamma_{{\underline{x}}}\kappa(x;\theta_{{\underline{x}}})\,\mathrm{d}x$, with $(\theta_{{\underline{x}}},\gamma_{{\underline{x}}})\sim f_{{\underline{x}}}(\mathrm{d}\theta\,\mathrm{d}\gamma)$ and $f_{{\underline{x}}}$ is defined in the statement of the present lemma. Consequently, the reduced Palm version $\left(\Phi_{\zeta_{{\underline{x}}}}\right)^{!}_{y}$ is a Cox process directed by $\mu_{y}$. We are then left with determining the distribution of $\mu_{y}$, which we derive by computing its Laplace functional $\mathcal{L}_{\mu_{y}}(f)$, for any measurable function $f:\mathbb{X}\to\mathbb{R}_{+}$. To this end, rely on (Baccelli et al., 2020, Proposition 3.2.1), which states that, for any measurable functions $f,g:\mathbb{X}\to\mathbb{R}_{+}$, the Palm distribution of a random measure $\mu$ satisfies
$\frac{\partial}{\partial t}\mathcal{L}_{\mu}(f+tg)\big|_{t=0}=-E\left[\mu(g)e^{-\mu(f)}\right]=-\int_{\mathbb{X}}g(y)\mathcal{L}_{\mu_{y}}(f)M_{\mu}(\mathrm{d}y),$ (S33)
where $\mu(f):=\int_{\mathbb{X}}f(x)\mu(\mathrm{d}x)$. Therefore, since $\mu(\mathrm{d}x)=\gamma_{{\underline{x}}}\kappa(x;\theta_{{\underline{x}}})\,\mathrm{d}x$, the term $E\left[\mu(g)e^{-\mu(f)}\right]$ writes as
$E\left[\mu(g)e^{-\mu(f)}\right]$ $=E\left[\int_{\mathbb{X}}g(y)\gamma_{\underline{x}}\kappa(y;\theta_{\underline{x}})\mathrm{d}y\cdot\exp\left\{-\int_{\mathbb{X}}f(z)\gamma_{\underline{x}}\kappa(z;\theta_{\underline{x}})\mathrm{d}z\right\}\right]$
$=\int_{\mathbb{X}\times\mathbb{R}_{+}}\int_{\mathbb{X}}g(y)\gamma\kappa(y;\theta)\mathrm{d}y\cdot\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\cdot f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)$
$=\int_{\mathbb{X}}g(y)\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\gamma\kappa(y;\theta)f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\cdot\mathrm{d}y.$ (S34)
Observing that $M_{\mu}(\mathrm{d}y)=E[\gamma_{\underline{x}}\kappa(y;\theta_{\underline{x}})\mathrm{d}y]=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(y;\theta)f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\,\mathrm{d}y$, (S34) can be expressed as
$E\left[\mu(g)e^{-\mu(f)}\right]=\int_{\mathbb{X}}g(y)\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\frac{\gamma\kappa(y;\theta)f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)}{\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{\prime}\kappa(y;\theta^{\prime})f_{\underline{x}}(\mathrm{d}\theta^{\prime}\,\mathrm{d}\gamma^{\prime})}\cdot M_{\mu}(\mathrm{d}y)$
@@
$\mathcal{L}_{\mu_{y}}(f)=\int_{\mathbb{X}\times\mathbb{R}_{+}}\exp\left\{-\int_{\mathbb{X}}f(z)\gamma\kappa(z;\theta)\mathrm{d}z\right\}\frac{\gamma\kappa(y;\theta)f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)}{\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{\prime}\kappa(y;\theta^{\prime})f_{\underline{x}}(\mathrm{d}\theta^{\prime}\,\mathrm{d}\gamma^{\prime})},$
@@
$(\theta^{*},\gamma^{*})\sim f_{*}(\mathrm{d}\theta\,\mathrm{d}\gamma)\propto\gamma\kappa(y;\theta)f_{\underline{x}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\propto f_{(\underline{x},y)}(\mathrm{d}\theta\,\mathrm{d}\gamma).$
@@
$\left(\Phi_{\zeta_{{\underline{x}}}}\right)^{!}_{y}\,|\,(\theta^{*},\gamma^{*})\sim\textsc{pp}(\gamma^{*}\kappa(x;\theta^{*})\,\mathrm{d}x),\qquad(\theta^{*},\gamma^{*})\sim f_{(\underline{x},y)}(\mathrm{d}\theta\,\mathrm{d}\gamma),$
which is equivalent to say $\left(\Phi_{\zeta_{{\underline{x}}}}\right)^{!}_{y}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\Phi_{\zeta_{({\underline{x}},y)}}$, and the thesis is proved.
Proof of Lemma˜S3
@@
$M^{(k)}_{\Phi}(B_{1},\ldots,B_{k})$ $=E\left[\sum_{X_{1},\ldots,X_{k}\in\Phi}^{\neq}\prod_{j=1}^{k}{1}_{B_{j}}(X_{j})\right]=E\left[E\left[\sum_{X_{1},\ldots,X_{k}\in\Phi}^{\neq}\prod_{j=1}^{k}{1}_{B_{j}}(X_{j})\,|\,\Lambda\right]\right]$
$=E\left[M^{(k)}_{\Phi\,|\,\Lambda}(B_{1},\ldots,B_{k})\right]=E\left[\prod_{j=1}^{k}M_{\Phi\,|\,\Lambda}(B_{j})\right]$
$=E\left[\prod_{j=1}^{k}\int_{B_{j}}\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(x_{j};\theta)\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}x_{j}\right]$
$=E\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{k}}\prod_{j=1}^{k}\gamma_{j}\int_{B_{j}}\kappa(x_{j};\theta_{j})\mathrm{d}x_{j}\Lambda^{k}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma})\right]$
$=\int_{(\mathbb{X}\times\mathbb{R}_{+})^{k}}\prod_{j=1}^{k}\gamma_{j}\int_{B_{j}}\kappa(x_{j};\theta_{j})\mathrm{d}x_{j}M_{\Lambda}^{k}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma}),$ (S35)
@@
Moreover, from (Baccelli et al., 2020, Lemma 14.E.4), the following holds true for any point process $\xi$ on a Polish space $\mathbb{Y}$ (embedded with the Borel $\sigma$-algebra $\mathscr{Y}$), for any $A_{1},\ldots,A_{k}\in\mathscr{Y}$,
@@
where $k_{h}$ is the cardinality of set $J_{h}$. For $\underline{x}_{J_{h}}=(x_{m}:m\in J_{h})$, define
$\eta(\underline{x}_{J_{h}})=\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma^{k_{h}}\prod_{m\in J_{h}}\kappa(x_{m};\theta)M_{\Lambda}(\mathrm{d}\theta\,\mathrm{d}\gamma).$